Water at 15 degrees Celcius (surface tension is 0.0742 N/m) stands in a clean glass tube of 1.5-mm diameter at a height of 45 mm. What is the true static height?
Answer #1

water at 150c Surface Tension (6) - 0.0 742 Nm Speefie weight (X) = 9798 N/m3 Diameter (D = 1.5 mm 21.5*102m height (H) = 45m